# Graph Neural Network based Log Anomaly Detection and Explanation

### v1

@@ Abstract.
Event logs are widely used to record the status of high-tech systems, making log anomaly detection important for monitoring those systems. Most existing log anomaly detection methods take a log event count matrix or log event sequences as input, exploiting quantitative and/or sequential relationships between log events to detect anomalies. Unfortunately, only considering quantitative or sequential relationships may result in many false positives and/or false negatives. To alleviate this problem, we propose a graph-based method for unsupervised log anomaly detection, dubbed Logs2Graphs, which first converts event logs into attributed, directed, and weighted graphs, and then leverages graph neural networks to perform graph-level anomaly detection. Specifically, we introduce One-Class Digraph Inception Convolutional Networks, abbreviated as OCDiGCN, a novel graph neural network model for detecting graph-level anomalies in a collection of attributed, directed, and weighted graphs. By coupling the graph representation and anomaly detection steps, OCDiGCN can learn a representation that is especially suited for anomaly detection, resulting in a high detection accuracy. Importantly, for each identified anomaly, we additionally provide a small subset of nodes that play a crucial role in OCDiGCN’s prediction as explanations, which can offer valuable cues for subsequent root cause diagnosis. Experiments on five benchmark datasets show that Logs2Graphs performs at least on par state-of-the-art log anomaly detection methods on simple datasets while largely outperforming state-of-the-art log anomaly detection methods on complicated datasets.
@@
Log Analysis, Log Anomaly Detection, Graph Neural Networks
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: xxxx; xx xxxx; xx, xxccs: Software and its engineering Maintaining software
1. Introduction
Modern high-tech systems, such as cloud computers or lithography machines, typically consist of a large number of components. Over time these systems have become larger and more complex, making manual system operation and maintenance hard or even infeasible. Therefore, automated system operation and maintenance is highly desirable. To achieve this, system logs are universally used to record system states and important events. By analysing these logs, faults and potential risks can be identified, and remedial actions may be taken to prevent severe problems. System logs are usually semi-structured texts though, and identifying anomalies through log anomaly detection is often challenging.
Since both industry and academia show great interest in identifying anomalies from logs, a plethora of log anomaly detection methods have been proposed. Existing log anomaly detection methods can be roughly divided into three categories: quantitative-based, sequence-based, and graph-based methods. Specifically, quantitative-based methods, such as OCSVM (Schölkopf et al., 2001) and PCA (Xu et al., 2009), utilise a log event count matrix to detect anomalies, and are therefore unable to capture semantic information of and sequential information between log events. Meanwhile, sequence-based methods, including DeepLog (Du et al., 2017) and LogAnomaly (Meng et al., 2019), aim to detect anomalies by taking sequential (and sometimes semantic) information into account. They cannot consider the full structure among log events though. In contrast, graph-based methods, such as GLAD-PAW (Wan et al., 2021), convert logs to graphs and exploit semantic information as well as the structure among log events, exhibiting the following three advantages over the former two categories of methods (Jia et al., 2017): 1) they are able to identify problems for which the structure among events is crucial, such as performance degradation; 2) they are capable of providing contextual log messages corresponding to the identified problems; and 3) they can provide the ‘normal’ operation process in the form of a graph, helping end-users find root-causes and take remedial actions. Existing graph-based methods, such as GLAD-PAW, typically transform log events into undirected graphs though, which may fail to capture important information on the order among log events. Moreover, most existing graph-based methods perform graph representation and anomaly detection separately, leading to suboptimal detection accuracy.
Moreover, as highlighted by Li et al. (Li et al., 2022), with the growing adoption of anomaly detection algorithms in safety-critical domains, there is a rising demand for providing explanations for the decisions made within those domains. This requirement, driven by ethical considerations and regulatory mandates, underscores the importance of accountability and transparency in such contexts. Moreover, in practical applications, the attainment of precise anomaly explanations contributes largely to the timely isolation and diagnosis of anomalies, which can mitigate the impact of anomalies by facilitating early intervention (Yang et al., 2021b). However, to our knowledge, most existing log anomaly detection methods focus exclusively on detection performance without giving any explanations.
To overcome these limitations, we propose Logs2Graphs, a graph-based unsupervised log anomaly detection approach by designing a novel one-class graph neural network. Specifically, Logs2Graphs first utilises off-the-shelf methods to learn a semantic embedding for each log event, and then assigns log messages to different groups. Second, Logs2Graphs converts each group of log messages into an attributed, directed, and weighted graph, with each node representing a log event, the node attributes containing its semantic embedding, a directed edge representing how an event is followed by another event, and the corresponding edge weight indicating the number of times the events follow each other. Third, by coupling the graph representation learning and anomaly detection objectives, we introduce One-Class Digraph Inception Convolutional Networks (OCDiGCN) as a novel method to detect anomalous graphs from a set of graphs. As a result, Logs2Graphs leverages the rich and expressive power of attributed, directed and edge-weighted graphs to represent logs, followed by using graph neural networks to effectively detect graph-level anomalies, taking into account both semantic information of log events and structure information (including sequential information as a special case) among log events. Importantly, by decomposing the anomaly score of a graph into individual nodes and visualizing these nodes based on their contributions, we provide straightforward and understandable explanations for identified anomalies.
Overall, our contributions can be summarised as follows: (1) We introduce Logs2Graphs, which formalises log anomaly detection as a graph-level anomaly detection problem and represents log sequences as directed graphs to capture more structure information than previous approaches; (2) We introduce OCDiGCN, the first end-to-end unsupervised graph-level anomaly detection method for attributed, directed and edge-weighted graphs. By coupling the graph representation and anomaly detection objectives, we improve the potential for accurate anomaly detection over existing approaches; (3) For each detected anomaly, we identify important nodes as explanations, offering valuable cues for subsequent root cause diagnosis; (4) We empirically compare our approach to eight state-of-the-art log anomaly detection methods on five benchmark datasets, showing that Logs2Graphs performs at least on par and often better than its competitors.
The reminder of this paper is organised as follows. Section 2 revisits related work, after which Section 3 formalises the problem. Section 4 describes Digraph Inception Convolutional Networks (Tong et al., 2020), which are used for Logs2Graphs in Section 5. We then evaluate Logs2Graphs in Section 6 and conclude in Section 7.
2. Related Work
Graph-based log anomaly detection methods usually comprise five steps: log parsing, log grouping, graph construction, graph representation learning, and anomaly detection. In this paper we focus on graph representation learning, log anomaly detection and explanation, thus only revisiting related work in these fields.
2.1. Graph Representation Learning
Graph-level representation learning methods, such as GIN (Xu et al., 2018) and Graph2Vec (Narayanan et al., 2017), are able to learn a mapping from graphs to vectors. Further, graph kernel methods, including Weisfeiler-Lehman (WL) (Shervashidze et al., 2011) and Propagation Kernels (PK) (Neumann et al., 2016), can directly provide pairwise distances between graphs. Both types of methods can be combined with off-the-shelf anomaly detectors, such as OCSVM (Schölkopf et al., 2001) and iForest (Liu et al., 2008), to perform graph-level anomaly detection.
To improve on these naïve approaches, efforts have been made to develop graph representation learning methods especially for anomaly detection. For instance, OCGIN (Zhao and Akoglu, 2021) and GLAM (Zhao et al., 2022) combine the GIN (Xu et al., 2018) representation learning objective with the SVDD objective (Tax and Duin, 2004) to perform graph-level representation learning and anomaly detection in an end-to-end manner. GLocalKD (Ma et al., 2022) performs random distillation of graph and node representations to learn ‘normal’ graph patterns. Further, OCGTL (Qiu et al., 2022) combines neural transformation learning and one-class classification to learn graph representations for anomaly detection. Although these methods are unsupervised or semi-supervised, they can only deal with attributed, undirected, and unweighted graphs.
iGAD (Zhang et al., 2022b) considers graph-level anomaly detection as a graph classification problem and combines attribute-aware graph convolution and substructure-aware deep random walks to learn graph representations. However, iGAD is a supervised method, and can only handle attributed, undirected, and unweighted graphs. CODEtect (Nguyen et al., 2020) takes a pattern-based modelling approach using the minimum description length (MDL) principle and identifies anomalous graphs based on motifs. CODEtect can (only) deal with labelled, directed, and edge-weighted graphs, but is computationally very expensive. To our knowledge, we introduce the first unsupervised method for graph-level anomaly detection that can handle attributed, directed and edge-weighted graphs.
2.2. Log Anomaly Detection and Explanation
Log anomaly detection methods can be roughly subdivided into three categories: 1) traditional, ‘shallow’ methods, such as principal component analysis (PCA) (Xu et al., 2009), one-class SVM (OCSVM) (Schölkopf et al., 2001), isolation forest (iForest) (Liu et al., 2008) and histogram-based outlier score (HBOS) (Goldstein and Dengel, 2012), which take a log event count matrix as input and analyse quantitative relationships; 2) deep learning based methods, such as DeepLog (Du et al., 2017), LogAnomaly (Meng et al., 2019), and AutoEncoder (Farzad and Gulliver, 2020), which employ sequences of log events (and sometimes their semantic embeddings) as input, analysing sequential information and possibly semantic information of log events to identify anomalies; and 3) graph-based methods, such as TCFG (Jia et al., 2017) and GLAD-PAW (Wan et al., 2021), which first convert logs into graphs and then perform graph-level anomaly detection.
To our knowledge, only a few works (Wan et al., 2021; Zhang et al., 2022a; Xie et al., 2022) have capitalised on the powerful learning capabilities of graph neural networks for log anomaly detection. GLAD-PAW (Wan et al., 2021) first transforms logs into attributed and undirected graphs and then uses a Position Aware Weighted Graph Attention Network to identify anomalies. However, converting logs into undirected graphs may result in loss of important sequential information. Further, DeepTraLog (Zhang et al., 2022a) combines traces and logs to generate a so-called Trace Event Graph, which is an attributed and directed graph. On this basis, they train a Gated Graph Neural Networks based Deep Support Vector Data Description model to identify anomalies. However, their approach requires the availability of both traces and logs, and is unable to handle edge weights. In contrast, like LogGD (Xie et al., 2022), our proposed Logs2Graphs approach is applicable to generic logs by converting logs into attributed, directed, and edge-weighted graphs. However, LogGD is a supervised method that requires fully labelled training data, which is usually impractical and even impossible. In contrast, our proposed algorithm OCDiGCN is the first unsupervised graph-level anomaly detection method for attributed, directed, and edge-weighted graphs.
Although anomaly explanation has received much attention in traditional anomaly detection (Li et al., 2022), only a few studies (Yang et al., 2021a) considered log anomaly explanation. Specifically, PLELog (Yang et al., 2021a) offers explanations by quantifying the significance of individual log events within an anomalous log sequence, thereby facilitating improved identification of relevant log events by operators. Similarly, our method provides straightforward explanations for anomalous log groups by identifying and visualising a small subset of important nodes.
3. Problem Statement
Before we state the log anomaly detection problem, we first introduce the necessary notations and definitions regarding event logs and graphs.
Event logs. Logs are used to record system status and important events, and are usually collected and stored centrally as log files. A log file typically consists of many log messages. Each log message is composed of three components: a timestamp, an event type (log event or log template), and additional information (log parameter). Log parsers are used to extract log events from log messages.
Further, log messages can be grouped into log groups (a.k.a. log sequences) using certain criteria. Specifically, if a log identifier is available for each log message, one can group log messages based on such identifiers. Otherwise, one can use a fixed or sliding window to group log messages. The window size can be determined according to timestamp or the number of observations. Besides, counting the occurrences of each log event within a log group results in a event count vector. Consequently, for a log file consisting of many log groups, one can obtain a event count matrix. The process of generating an event count matrix (or other feature matrix) is known as feature extraction. Extracted features are often used as input to an anomaly detection algorithm to identify log anomalies, i.e., log messages or log groups that deviate from what is considered ‘normal’.
Graphs. We consider an attributed, directed, and edge-weighted graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\textbf{X},\textbf{Y})$, where $\mathcal{V}=\{v_{1},...,v_{|\mathcal{V}|}\}$ denotes the set of nodes and =$\mathcal{E}=\{e_{1},...,e_{|\mathcal{E}|}\}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges between nodes. If $(v_{i},v_{j})\in\mathcal{E}$, then there is an edge from node $v_{i}$ to node $v_{j}$. Moreover, $\textbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the node attribute matrix, with the $i$-th row representing the attributes of node $v_{i}$, and $d$ is the number of attributes. Besides, $\textbf{Y}\in\mathbb{N}^{|\mathcal{E}|\times|\mathcal{E}|}$ is the edge-weight matrix, where $\textbf{Y}_{ij}$ represents the weight of the edge from node $v_{i}$ to $v_{j}$.
@@ 3. Problem Statement
3.1. Graph-based Log Anomaly Detection

### v3

# Graph Neural Networks based Log Anomaly Detection and Explanation

@@ Abstract
Event logs are widely used to record the status of high-tech systems, making log anomaly detection important for monitoring those systems. Most existing log anomaly detection methods take a log event count matrix or log event sequences as input, exploiting quantitative and/or sequential relationships between log events to detect anomalies. However, only considering quantitative or sequential relationships may result in low detection accuracy. To alleviate this problem, we propose a graph-based method for unsupervised log anomaly detection, dubbed Logs2Graphs, which first converts event logs into attributed, directed, and weighted graphs, and then leverages graph neural networks to perform graph-level anomaly detection. Specifically, we introduce One-Class Digraph Inception Convolutional Networks, abbreviated as OCDiGCN, a novel graph neural network model for detecting graph-level anomalies in a collection of attributed, directed, and weighted graphs. By integrating graph representation and anomaly detection, OCDiGCN learns a specialized representation that leads to high detection accuracy. Crucially, we furnish a concise set of nodes pivotal in OCDiGCN’s prediction as explanations for each detected anomaly, offering valuable insights for subsequent root cause analysis. Experiments on five benchmark datasets show that Logs2Graphs exhibits comparable or superior performance when compared to state-of-the-art log anomaly detection methods.
keywords:
@@
Log Analysis, Log Anomaly Detection, Graph Neural Networks, Smart Manufacturing
@@
[inst1]organization=LIACS, Leiden University,city=Leiden, country=the Netherlands
1 Introduction
Modern high-tech systems, such as cloud computers or lithography machines, typically consist of a large number of components. Over time these systems have become larger and more complex, making manual system operation and maintenance hard or even infeasible Li and van Leeuwen (2022). Therefore, automated system operation and maintenance is highly desirable. To achieve this, system logs are universally used to record system states and important events. By analysing these logs, faults and potential risks can be identified, and remedial actions may be taken to prevent severe problems. System logs are usually semi-structured texts though, and identifying anomalies through log anomaly detection is often challenging.
Since both industry and academia show great interest in identifying anomalies from logs, a plethora of log anomaly detection methods have been proposed. Existing log anomaly detection methods can be roughly divided into three categories: quantitative-based, sequence-based, and graph-based methods. Specifically, quantitative-based methods, such as OCSVM Schölkopf et al. (2001) and PCA Xu et al. (2009), utilise a log event count matrix to detect anomalies, and are therefore unable to capture semantic information of and sequential information between log events. Meanwhile, sequence-based methods, including DeepLog Du et al. (2017) and LogAnomaly Meng et al. (2019), aim to detect anomalies by taking sequential (and sometimes semantic) information into account. They cannot consider the full structure among log events though. In contrast, graph-based methods, such as GLAD-PAW Wan et al. (2021) and GLAD Li et al. (2023a), convert logs to graphs and exploit semantic information as well as the structure among log events, exhibiting the following three advantages over the former two categories of methods Jia et al. (2017): 1) they are able to identify problems for which the structure among events is crucial, such as performance degradation; 2) they are capable of providing contextual log messages corresponding to the identified problems; and 3) they can provide the ‘normal’ operation process in the form of a graph, helping end-users find root-causes and take remedial actions. However, graph-based methods like GLAD transform log events into undirected graphs though, which may fail to capture important information on the order among log events. Moreover, most existing graph-based methods perform graph representation and anomaly detection separately, leading to suboptimal detection accuracy.
Moreover, as highlighted by Li et al. Li et al. (2023b), with the growing adoption of anomaly detection algorithms in safety-critical domains, there is a rising demand for providing explanations for the decisions made within those domains. This requirement, driven by ethical considerations and regulatory mandates, underscores the importance of accountability and transparency in such contexts. Moreover, in practical applications, the attainment of precise anomaly explanations contributes to the timely isolation and diagnosis of anomalies, which can mitigate the impact of anomalies by facilitating early intervention Yang et al. (2021b). However, to our knowledge, most existing log anomaly detection methods focus exclusively on accurate detection without giving explanations.
To overcome these limitations, we propose Logs2Graphs, a graph-based unsupervised log anomaly detection approach for which we design a novel one-class graph neural network. Specifically, Logs2Graphs first utilises off-the-shelf methods to learn a semantic embedding for each log event, and then assigns log messages to different groups. Second, Logs2Graphs converts each group of log messages into an attributed, directed, and weighted graph, with each node representing a log event, the node attributes containing its semantic embedding, directed edges representing how an event is followed by other events, and the corresponding edge weights indicating the number of times the events follow each other. Third, by coupling the graph representation learning and anomaly detection objectives, we introduce One-Class Digraph Inception Convolutional Networks (OCDiGCN) as a novel method to detect anomalous graphs from a set of graphs. As a result, Logs2Graphs leverages the rich and expressive power of attributed, directed, and edge-weighted graphs to represent logs, followed by using graph neural networks to effectively detect graph-level anomalies, taking into account both semantic information of log events and structure information (including sequential information as a special case) among log events. Importantly, by decomposing the anomaly score of a graph into individual nodes and visualizing these nodes based on their contributions, we provide understandable explanations for identified anomalies.
Overall, our contributions can be summarised as follows: (1) We introduce Logs2Graphs, which formalises log anomaly detection as a graph-level anomaly detection problem and represents log sequences as directed graphs to capture more structure information than previous approaches; (2) We introduce OCDiGCN, a general end-to-end unsupervised graph-level anomaly detection method for attributed, directed and edge-weighted graphs. By coupling the graph representation and anomaly detection objectives, we improve the potential for accurate anomaly detection over existing approaches; (3) For each detected anomaly, we identify important nodes as explanations, offering cues for subsequent root cause diagnosis; (4) We empirically compare our approach to eight state-of-the-art log anomaly detection methods on five benchmark datasets, showing that Logs2Graphs performs at least on par with and often better than its competitors.
2 Related Work
Graph-based log anomaly detection methods usually comprise five steps: log parsing, log grouping, graph construction, graph representation learning, and anomaly detection. In this paper we focus on graph representation learning, log anomaly detection, and explanation, thus only revisiting related work in these fields.
2.1 Graph Representation Learning
Graph-level representation learning methods, such as GIN Xu et al. (2018) and Graph2Vec Narayanan et al. (2017), are able to learn a mapping from graphs to vectors. Further, graph kernel methods, including Weisfeiler-Lehman (WL) Shervashidze et al. (2011) and Propagation Kernels (PK) Neumann et al. (2016), can directly provide pairwise distances between graphs. Both types of methods can be combined with off-the-shelf anomaly detectors, such as OCSVM Schölkopf et al. (2001) and iForest Liu et al. (2008), to perform graph-level anomaly detection.
To improve on these naïve approaches, efforts have been made to develop graph representation learning methods especially for anomaly detection. For instance, OCGIN Zhao and Akoglu (2021) and GLAM Zhao et al. (2022) combine the GIN Xu et al. (2018) representation learning objective with the SVDD objective Tax and Duin (2004) to perform graph-level representation learning and anomaly detection in an end-to-end manner. GLocalKD Ma et al. (2022) performs random distillation of graph and node representations to learn ‘normal’ graph patterns. Further, OCGTL Qiu et al. (2022) combines neural transformation learning and one-class classification to learn graph representations for anomaly detection. Although these methods are unsupervised or semi-supervised, they can only deal with attributed, undirected, and unweighted graphs.
iGAD Zhang et al. (2022b) considers graph-level anomaly detection as a graph classification problem and combines attribute-aware graph convolution and substructure-aware deep random walks to learn graph representations. However, iGAD is a supervised method, and can only handle attributed, undirected, and unweighted graphs. CODEtect Nguyen et al. (2023) takes a pattern-based modelling approach using the minimum description length principle and identifies anomalous graphs based on motifs. CODEtect can (only) deal with labelled, directed, and edge-weighted graphs, but is computationally very expensive. In contrast, we introduce a general unsupervised method for graph-level anomaly detection that can handle attributed, directed and edge-weighted graphs.
2.2 Log Anomaly Detection and Explanation
Log anomaly detection methods can be roughly divided into: 1) traditional, ‘shallow’ methods, such as principal component analysis (PCA) Xu et al. (2009), one-class SVM (OCSVM) Schölkopf et al. (2001), isolation forest (iForest) Liu et al. (2008), and histogram-based outlier score (HBOS) Goldstein and Dengel (2012), which take a log event count matrix as input and analyse quantitative relationships; 2) deep learning based methods, such as DeepLog Du et al. (2017), LogAnomaly Meng et al. (2019), and AutoEncoder Farzad and Gulliver (2020), which employ sequences of log events (and sometimes their semantic embeddings) as input, analysing sequential information and possibly semantic information of log events to identify anomalies; and 3) graph-based methods, such as TCFG Jia et al. (2017) and GLAD-PAW Wan et al. (2021), which first convert logs into graphs and then perform graph-level anomaly detection.
To our knowledge, only a few works Wan et al. (2021); Zhang et al. (2022a); Xie et al. (2022); Li et al. (2023a) have capitalised on the powerful learning capabilities of graph neural networks for log anomaly detection. GLAD Li et al. (2023a) transforms logs into undirected, weighted, and attributed heterogeneous graph, and propose a temporal-attentive graph edge anomaly detection model to detect anomalous relations. However, converting logs into undirected graphs may result in loss of important sequential information. Further, DeepTraLog Zhang et al. (2022a) combines traces and logs to generate a so-called Trace Event Graph, which is an attributed and directed graph. On this basis, they train a Gated Graph Neural Networks based Deep Support Vector Data Description model to identify anomalies. However, their approach requires the availability of both traces and logs, and is unable to handle edge weights. In contrast, like LogGD Xie et al. (2022) and GLAD-PAW Wan et al. (2021), our proposed Logs2Graphs approach is applicable to generic logs by converting logs into attributed, directed, and edge-weighted graphs. However, LogGD is a supervised method that requires fully labelled training data, which is usually impractical and even impossible. Moreover, LogGD and GLAD-PAW Wan et al. (2021) are not capable of providing explanations for identified anomalies. In contrast, our proposed algorithm OCDiGCN is a general unsupervised graph-level anomaly detection method for attributed, directed, and edge-weighted graphs, able to provide explanations for identified anomalies.
Although anomaly explanation has received much attention in traditional anomaly detection Li et al. (2023b); Li and van Leeuwen (2023), only a few studies Yang et al. (2021a) considered log anomaly explanation. Specifically, PLELog Yang et al. (2021a) offers explanations by quantifying the significance of individual log events within an anomalous log sequence, thereby facilitating improved identification of relevant log events by operators. Similarly, our method provides explanations for anomalous log groups by identifying and visualising a small subset of important nodes.
3 Problem Statement
Before we state the log anomaly detection problem, we first introduce notation and definitions regarding event logs and graphs.
Event logs. Logs are used to record system status and important events, and are usually collected and stored centrally as log files. A log file typically consists of many log messages. Each log message is composed of three components: a timestamp, an event type (log event or log template), and additional information (log parameters). Log parsers are used to extract log events from log messages.
Further, log messages can be grouped into log groups (a.k.a. log sequences) using certain criteria. Specifically, if a log identifier is available for each log message, one can group log messages based on such identifiers. Otherwise, one can use a fixed or sliding window to group log messages. Besides, counting the occurrences of each log event within a log group yields an event count vector. Consequently, for a log file consisting of many log groups, one can obtain an event count matrix. The process of generating an event count matrix (or other feature matrix) is known as feature extraction. Extracted features are often used as input to an anomaly detection algorithm to identify log anomalies, i.e., log messages or log groups that deviate from what is considered ‘normal’.
Graphs. We consider an attributed, directed, and edge-weighted graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\textbf{X},\textbf{Y})$, where $\mathcal{V}=\{v_{1},...,v_{|\mathcal{V}|}\}$ denotes the set of nodes and =$\mathcal{E}=\{e_{1},...,e_{|\mathcal{E}|}\}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges. If $(v_{i},v_{j})\in\mathcal{E}$, then there is an edge from node $v_{i}$ to node $v_{j}$. Moreover, $\textbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the node attribute matrix, with the $i$-th row representing the attributes of node $v_{i}$, and $d$ is the number of attributes. Besides, $\textbf{Y}\in\mathbb{N}^{|\mathcal{E}|\times|\mathcal{E}|}$ is the edge-weight matrix, where $\textbf{Y}_{ij}$ represents the weight of the edge from node $v_{i}$ to $v_{j}$.
[…]
3.1 Graph-based Log Anomaly Detection